susan is 3 times as old as blake is now. In fifteen years, she will be 10 years older than blake. how old are they?

Answer by ikleyn(52790)

.

Read this sentence "In fifteen years, she will be 10 years older than Blake."


It has very deep meaning.


It means that NOT ONLY in 15 years Susan will be 10 years older than Blake.

It means that ALWAYS Susan will be (and is; and was) 10 years older than Blake.


In particular, right now she IS 10 years older than Blake.


In other words,  S = B + 10,

therefore, for their current age you have this equation

    3B = B + 10,


which implies

    3B - B = 10

    2B     = 10

     B     = 10/2 = 5.


ANSWER.  Blake is 5 years old; Susan is 3*5 = 15 years old.